# ELUTQ: Efficient LUT-Aware Quantization for Deploying Large Language Models on Edge Devices

### v2

@@ Appendix A Gradient-Based Search for Hierarchical Linear Quantization
Appendix B Compression Ratios and Actual Memory Footprint of Quantized Models
@@
Appendix C Compared with Non-Uniform and Quantization-Aware Training Methods
In this section, we further compare our results with several non-uniform quantization that adopt codebooks and QAT methods. As shown in the table 16, HLQ-Finetuning exhibit a certain gap in model accuracy compared to these non-uniform approaches, which is expected. Methods such as Quip# (Tseng et al., 2024) and AQLM (Egiazarian et al., 2024) require a large calibration set (e.g., 4K training samples with a context length of 4096) and involve complex distillation procedures, leading to substantial computational and time costs during quantization. Meanwhile, when using the same training data, our method achieves comparable results to EfficientQAT (Chen et al., 2024c), however, EfficientQAT necessitates retraining the model weights, whereas our approach operates without any weight training, significantly reducing memory consumption, accelerating quantization, and mitigating the risk of overfitting.
Table 16: Evaluation of quantized LLaMA3.1 models for 2 bits. The table reports perplexity on WikiText2 and c4, as well as accuracy for zero-shot tasks. The Average column is the mean of 5 zero-shot task accuracies. EfficientQAT* indicates the version trained using the same dataset as HLQ-Finetuning.
@@ Appendix C Compared with Non-Uniform and Quantization-Aware Training Methods
Quip# 9.92 13.45 69.78 54.18 40.12 74.56 78.75 63.48
AQLM 9.56 12.98 70.92 54.92 40.84 74.48 78.12 63.85
EfficientQAT 10.12 14.95 64.58 49.89 36.13 68.34 74.56 58.71
EfficientQAT* 11.56 16.34 62.58 49.76 33.13 68.34 74.56 57.68
HLQ-Finetuning 11.24 15.87 62.72 49.52 32.91 69.02 74.78 57.79
LlaMA3.1-70B Baseline 2.93 6.92 80.51 66.36 60.41 86.99 82.37 75.33
Quip# 6.89 10.47 77.11 61.94 50.45 77.05 78.32 68.98
AQLM 6.38 9.73 77.56 62.87 50.21 78.26 79.15 69.61
EfficientQAT 6.92 10.56 69.27 59.95 49.04 78.84 79.22 67.26
EfficientQAT * 7.66 11.48 71.57 57.66 46.29 76.80 77.72 66.00
HLQ-Finetuning 7.48 11.24 72.38 58.86 46.52 77.16 77.54 66.50
Appendix D Experimental Configurations for Memory and Time Comparison.
@@
In this section, we provide the detailed experimental configurations used to measure the computational memory consumption (CPU and GPU) and quantization time of different quantization methods on LLaMA3.1–70B under 2-bit quantization. All the experiments is conducted on RTX 4090.
@@
  • GPTQ: Following the official implementation, we use 128 calibration samples with a context length of 2048 and perform layer-wise quantization.
@@
  • AWQ: According to the official codebase, we use 512 calibration samples from the default AWQ dataset, each with a context length not exceeding 512.
@@
  • OmniQuant: We use 128 calibration samples with a context length of 2048 and follow the official recommendation to iterate 20 epochs for each transformer block.
  • HLQ-GPTQ: We adopt the same configuration as GPTQ, using 128 calibration samples with a context length of 2048 for layer-wise quantization.
@@
  • Quip#: We use 4K calibration samples with a context length of 4096 and conduct both layer-wise quantization.
@@
  • AQLM: We follow the same configuration as Quip#, using 4K calibration samples with a context length of 4096, and perform block-wise quantization.
  • DB-LLM: We use a self-collected dataset of 20K samples for calibration.
@@
  • EfficientQAT: We use 1K calibration samples with a context length of 2048,iterate 2 epochs for each transformer block.
@@
  • HLQ-Finetuning: We use 1K calibration samples with a context length of 2048,iterate 2 epochs for each transformer block.
Appendix E More results on Devices
E.1 More Results on ARM-Based Devices
@@
E.2 Results on GPUs
Here, we provide the experimental results on GPUs. Our evaluation covers a range of hardware platforms, including high-performance server-grade chips (A100 and H100), a consumer-grade GPU (RTX 4090), and an edge-computing–oriented accelerator (Jetson AGX Orin). As baselines, we consider AWQ, which uses a uniform quantization format, and SqueezeLLM, which adopts a non-uniform quantization format. We convert our HLQ weights into the BCQ format via a simple linear transformation and perform inference using the AnyBCQ framework. The experiments evaluate the speedup of throughput for generating a sequence of length 1024 compared with fp16 using cu, and the results are presented in the table 18.
@@ 3 AWQ 34 54 OOM OOM
Appendix F System-level Optimizations for the Quantization Pipeline
F.1 Lazy Loading
In both GPTQ’s Hessian-based quantization and Efficient Finetuning’s block reconstruction, the quantization unit is a Transformer block. Quantizing a single block does not depend on other blocks, neither the already quantized nor the unquantized ones. Therefore, instead of loading the entire model into memory, we can load only one transformer block at a time from disk. Thanks to the sharded weight storage mechanism in the Transformers (Wolf et al., 2020) framework, we can leverage the weight mapping file (e.g., model.safetensors.index.json) to selectively load a specific block from disk for quantization. Once the quantization of a block on the GPU is completed, it is immediately offloaded to the CPU to reduce GPU memory usage.
F.2 Chunked Computation of HLQ Parameters
In Algorithm 1, when computing the initial quantization parameters of HLQ for a weight matrix, the optimization within each weight group is independent. For extremely large weight matrices, we can split the matrix into smaller chunks and perform parameter search independently for each chunk. As shown in the figure, for the largest matrix in Llama3.1-70B, the down_proj of shape [28672, 8192] performing HLQ parameter search without batching leads to a peak memory usage of approximately 30 GB for 2-bit quantization and 35 GB for 3-bit quantization. In contrast, when the search is executed in two chunks, the peak memory drops to about 22 GB for 2-bit and 24 GB for 3-bit quantization. This chunked computation strategy significantly reduces GPU memory consumption and achieves approximately linear memory scaling with respect to partition granularity.
F.3 Offloading Hidden States to Disk
F.4 Overhead of Disk Access
Using lazy loading and loading training data directly from disk requires disk access during the execution of quantization algorithms. This inevitably introduces additional memory access overhead. Even with NVMe-based communication protocols, disk bandwidth remains far below that of the CPU. As shown in Table 19, we evaluate HLQ-GPTQ and HLQ-Finetuning on LLaMA3.1-8B, Qwen3-32B, and LLaMA3.1-70B. For HLQ-GPTQ, enabling disk-based loading substantially increases the quantization time, by approximately 30%–40% compared to in-memory loading. However, in the case of HLQ-Finetuning, the impact of disk access is relatively minor compared to the overall tuning time. Specifically, the tuning time increases by 25% for LLaMA3.1-8B, 12% for Qwen3-32B, and only about 5% for LLaMA3.1-70B. Therefore, when performing HLQ-Finetuning on very large models, the overhead caused by disk access can be considered negligible relative to the total tuning time.
Table 19: Comparison of tuning times with and without disk usage. All models are tuned under the same quantization configuration w2g128.
Method Model disk(✗) disk(✓) Overhead (%)
HLQ-GPTQ llama3.1-8B 0.25h 0.35h 40%
Qwen3-32B 1.35h 1.85h 37%
llama3.1-70B 3.00h 4.00h 33%
HLQ-Finetuning llama3.1-8B 2.00h 2.50h 25%
Qwen3-32B 8.25h 9.30h 12%
llama3.1-70B 19.00 20.00h 5%
Appendix G Full Results on Model Accuracy
Here, we present the full results of our method across various models. Specifically, Table 20 reports the perplexity on wikitext2, Table 21 reports the accuracy on the LLaMA series, while Table 22 shows the accuracy on the Qwen series models.
@@ Appendix G Full Results on Model Accuracy
AnyBCQ 2 128 2.37 - - 19.01 - - - -
HLQ-GPTQ 2 128 2.37 9.90 7.02 17.32 14.42 18.15 17.54 11.37
HLQ-Finetuning 2 128 2.37 8.05 6.75 11.24 8.43 17.87 14.92 10.31
DB-LLM 2 64 - 7.23 6.19 13.60 - - - -
HLQ-GPTQ 2 64 2.75 8.72 6.47 13.22 11.77 17.13 13.11 10.47
HLQ-Finetuning 2 64 2.75 7.07 6.07 11.22 7.06 15.09 12.24 9.72
@@
OmniQ 3 128 3.25 6.03 5.28 8.27 5.66 - - -
SqueezeLLM(0.45%) 3 - 3.25 5.96 5.23 - - - - -
GANQ 3 - 3.25 5.96 - 7.46 - - - -
@@
HLQ-GPTQ 2 128 59.27 38.96 22.18 40.70 60.66 44.35
HLQ-Finetuning 2 128 62.72 49.52 32.91 69.02 74.78 57.79
LLaMA3.1-70B Baseline 16 - 80.51 66.36 60.41 86.99 82.37 75.33
GPTQ 3 128 78.14 62.58 52.99 82.07 80.63 71.28

### v3

@@ Appendix D Design of High Performance CPU and GPU kernels
Appendix E Compression Ratios and Actual Memory Footprint of Quantized Models
@@
Appendix F Compared with Non-Uniform and Quantization-Aware Training Methods
In this section, we further compare our results with several codebook-based non-uniform quantization that adopt codebooks and QAT methods. As shown in the table 9, ELUTQ exhibit a certain gap in model accuracy compared to these non-uniform approaches, which is expected. Methods such as Quip# Tseng et al. [2024] and AQLM Egiazarian et al. [2024] require a large calibration set (e.g., 4K training samples with a context length of 4096) and involve complex distillation procedures, leading to substantial computational and time costs during quantization. Meanwhile, when using the same training data, our method achieves comparable results to EfficientQAT Chen et al. [2024b], however, EfficientQAT necessitates retraining the model weights, whereas our approach operates without any weight training, significantly reducing memory consumption, accelerating quantization, and mitigating the risk of overfitting.
Table 9: Evaluation of quantized LLaMA3.1 models for 2 bits. The table reports perplexity on WikiText2 and c4, as well as accuracy for zero-shot tasks. The Average column is the mean of 5 zero-shot task accuracies.
@@ Appendix F Compared with Non-Uniform and Quantization-Aware Training Methods
Quip# 11.81 14.24 65.78 51.18 38.12 70.56 76.75 60.48
AQLM 11.67 13.56 65.92 51.92 38.84 70.48 76.12 60.66
EfficientQAT 13.12 14.95 64.17 49.31 34.22 69.23 74.43 58.27
ELUTQ 12.38 15.08 62.72 49.52 32.91 69.02 74.78 57.79
LlaMA3.1-70B Baseline 2.93 6.92 80.72 66.58 60.62 86.99 82.77 75.52
Quip# 6.89 10.47 73.11 60.94 50.45 77.05 78.32 67.98
AQLM 6.38 9.73 73.56 61.87 50.21 78.26 79.15 68.61
EfficientQAT 7.35 10.56 69.27 59.95 49.04 78.84 79.22 67.26
ELUTQ 7.08 10.65 72.38 58.86 46.52 77.16 77.54 66.70
Appendix G More details of Efficiency Evaluation
G.1 Experimental Configurations for Memory and Time Comparison
@@
In this section, we provide the detailed experimental configurations used to measure the computational memory consumption (CPU and GPU) and quantization time of different quantization methods on LLaMA3.1–70B under 2-bit quantization. All the experiments is conducted on RTX 4090s.
@@
  • GPTQ: Following the official implementation, we use 128 calibration samples with a context length of 2048 and perform layer-wise quantization, using 2 RTX4090s.
@@
  • AWQ: According to the official codebase, we use 512 calibration samples from the default AWQ dataset, each with a context length not exceeding 512, using single RTX4090.
@@
  • OmniQuant: We use 128 calibration samples with a context length of 2048 and follow the official recommendation to iterate 20 epochs for each transformer block, using 2 RTX4090s.
@@
  • Quip#: We use 4K calibration samples with a context length of 4096 and conduct both layer-wise quantization, using 4 A100-80GB GPUs.
@@
  • AQLM: We follow the same configuration as Quip#, using 4K calibration samples with a context length of 4096, and perform block-wise quantization,using 4 A100-80GB GPUs.
  • DB-LLM: We use a self-collected dataset of 20K samples for calibration, using 4 A100-80GB GPUs.
@@
  • EfficientQAT: We use 4K calibration samples with a context length of 2048,iterate 2 epochs for each transformer block, using 2 RTX4090s in block-wise reconstruction and 2 RTX4090s in end to end tuning.
@@
  • ELUTQ: We use 4K calibration samples with a context length of 2048,iterate 2 epochs for each transformer block, using single RTX4090 in block-wise reconstruction and 2 RTX4090s in end to end tuning.
G.2 Parameter Efficiency
ELUTQ updates only a small subset of parameters in both stages, making it highly parameter-efficient. For LLaMA3.1-8B, we conduct experiments comparing the number of trainable parameters required for fine-tuning a single Transformer block and the entire model under different strategies, as shown in table 10. Our method is parameter-efficient.
Table 10: Comparison of the number of trainable parameters required by different fine-tuning strategies for w2g128 LLaMA3.1-8B.
Strategy Block-level Model-level
W 218M 7G
s,z,round 221M 7.1G
clipping 3M 0.1G
round 218M 7G
s,z,W 221M 7.1G
ELUTQ(ours) 5M 0.15G
Appendix H Full Results of Inference Performance on Edge Devices
H.1 Evaluation on CPUs
Table 11: Comparison of GEMV latency (ms) between uniform quantization and HLQ formats using the T-MAC kernel.
Chip Kernel $11008\times 4096$ $4096\times 32000$
T-MAC-HLQ 5.72(± 0.11)5.72\,($\pm\,$0.11) 22.23(± 0.16)22.23\,($\pm\,$0.16)
T-MAC-HLQ 1.86(± 0.04)1.86\,($\pm\,$0.04) 4.49(± 0.05)4.49\,($\pm\,$0.05)
Method W G Mem(GB) Speedup
[…]
Table 13: Comparison of prefill and decode throughput (tokens/s) between ELUTQ and llama.cpp on the Apple M2 chip. All experiments are conducted with 4 threads and a batch size of 1.
Model Framework Wbits BPW Input,Output Prefill Decode
llama2-7B llama.cpp Q3_K_M 3.91 128,128 19.87 15.65
256,128 19.75 15.53
Q3_K_S 3.50 128,128 17.60 15.44
256,128 17.43 15.00
Q2_K_S 2.50 128,128 14.82 13.35
256,128 14.54 13.16
ELUTQ W3(g128) 3.50 128,128 20.60 16.63
256,128 21.54 16.21
W2(g128) 2.37 128,128 25.06 20.48
256,128 25.94 20.25
@@ H.1 Evaluation on CPUs
H.2 Evaluation on GPUs
Here, we provide the experimental results on GPUs. Our evaluation covers a range of hardware platforms, including high-performance server-grade chips (A100 and H100), consumer-grade GPUs (RTX 4090 and RTX 3090). As baselines, we consider AWQ, which uses a uniform quantization format, and SqueezeLLM, which adopts a non-uniform quantization format. We convert our HLQ weights into the BCQ format via a simple linear transformation and perform inference using the AnyBCQ framework. The results are presented in the table 15.
@@ 3 AWQ 34 54 OOM OOM
Appendix I Full Results on Model Accuracy
Here, we present the full results of our method across various models. Specifically, Table 16 reports the perplexity on wikitext2, Table 17 reports the accuracy on the LLaMA series models, while Table 18 shows the accuracy on the Qwen series models.
@@ Appendix I Full Results on Model Accuracy
AnyBCQ 2 128 2.37 8.96 7.14 19.01 10.17 - - -
ELUTQ 2 128 2.37 8.05 6.75 12.38 7.08 17.87 14.92 10.31
DB-LLM 2 64 - 7.23 6.19 13.11 - - - -
ELUTQ 2 64 2.75 7.07 6.07 9.72 6.56 15.09 12.24 9.72
@@
OmniQ 3 128 3.25 6.03 5.28 8.27 5.66 11.71 9.86 9.04
SqueezeLLM(0.45%) 3 - 3.25 5.96 5.23 7.48 - - - -
GANQ 3 - 3.25 5.96 5.20 7.46 - - - -
@@
ELUTQ 2 128 62.72 49.52 32.91 69.02 74.78 57.79
LLaMA3.1-70B Baseline 16 - 80.72 66.58 60.62 86.99 82.77 75.52
GPTQ 3 128 77.94 62.76 52.59 81.74 80.50 71.08